OpenStudy (anonymous):

sqr root over sqr root x ^ 4

OpenStudy (anonymous):

first whats the square root of \[x ^{4}\]

OpenStudy (lgbasallote):

sq root over...?

Parth (parthkohli):

\[\large \sqrt{x^4} = {\left(x^4\right)^{1 \over 2}} \]

hartnn (hartnn):

remember \[\sqrt{a^2}=a\] and \(a^4=a^2.a^2\)

Parth (parthkohli):

Now use the property \((x^a)^b = x^{ab}\)

OpenStudy (lgbasallote):

am i really the only one seeing this sq root over?

OpenStudy (lgbasallote):

\[\huge \frac{\sqrt {}}{\sqrt{x^4}}\]

OpenStudy (anonymous):

\[\sqrt{\sqrt{x}} ^4\]

OpenStudy (lgbasallote):

oh..that

OpenStudy (lgbasallote):

well that's a totally different story then

OpenStudy (anonymous):

first the try to solve the first square root. i.e. \[\sqrt{x ^{4}}\]

OpenStudy (anonymous):

\[x ^{4 \times \frac{ 1 }{ 2 }}\]

OpenStudy (anonymous):

um okay, then what?

OpenStudy (anonymous):

then again apply the root for the result... whats the result?

OpenStudy (anonymous):

x^2 ?

OpenStudy (anonymous):

now 

OpenStudy (anonymous):

i.e. \[x ^{2 \times \frac{ 1 }{ 2 }}\]

OpenStudy (anonymous):

x^ 1 ?

OpenStudy (anonymous):

yeah..... i.e. x.

OpenStudy (anonymous):

would that be the asnwer then ?

OpenStudy (anonymous):

yeah!

OpenStudy (anonymous):

thanks @ :)

OpenStudy (anonymous):

you are welcomed....